Each of the letters D E T E R M I N E D is written on a card. YOu mix the cards thoroughly. What is the probability of selecting an E, then an M? The first card is replaced before selecting the second card.
Found 2 solutions by stanbon, Fombitz:
Answer by stanbon(75887)

Each of the letters D E T E R M I N E D is written on a card. YOu mix the cards thoroughly. What is the probability of selecting an E, then an M? The first card is replaced before selecting the second card.
----------
Prob (E) = 3/10
Prob (M) = 1/10
---------------
The draw results are independent because you replaced the 1st card drawn.
Therefore P(E then M) = (3/10)(1/10)= 3/100
=================
Cheers,
Stan H.

Answer by Fombitz(32388)

DETERMINED
2 D
3 E
1 T
1 R
1 M
1 I
1 N
10 letters total
Probability of selecting an E is 3/10.
Probability of selecting an M is 1/10.
Since the events are independent and the first card is replaced, the probability of E then M is then the product of the two probabilities.
P(E then M)=3/10*1/10=3/100.